• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Solvignon

    Log entry time 08:03:29 on July22,2006

    Entry number 174800

    keyword=Shift summary

    00:00 start the shift with no beam
    01:00 call MCC to have an update. They are working on Hall C optics.
    Request controlled access for Alexandre to check the retiming cable.
    01:30 Alexandre found out ROC 3 crashed.
    02;20 Bob is getting in the Hall to help Alexandre
    02:45 Bob exited the Hall and informed us that ROC 3 power supply is dead but we can still run with ROC 4 to test the crystals.
    02:55 Alexandre exited the Hall and we requested beam permit. Now the left arm DAQ should be run under fpp configuration.
    06:11 Vesda alarm beam way bleaking "urgent". Called MCC to report it. They are sending somebody to look at it.
    06:26 MCC called. They want to try to send 20uA. We moved the target to empty.
    06:33 Right arm Q1 at zero and doesn't come up by resetting from Counting house.
    06:35 Target vaccuum rising suddenly. Page Jian-Ping.
    06:42 MCC called to inform us that they had an incident around the compton chicane which caused a sudden lost of vaccuum.
    07:00 Target computer "froze".
    07:10 Jian-Ping said that target vaccuum problem is certainly related to the compton chicane vaccuum problem. Following the instructions from Jian-Ping, we rebooted the target computer. Now the target vaccuum is at 3.5E-7 which is good.
    07:40 MCC put Hall A in controlled access.
    08:00 Page Jack to replace ROC 3 power supply.